OpenStudy (anonymous):

Let f(x) = (3x^2 − 12) /(x^2 − 2x − 3) . Sketch a complete graph of f. Include all vertical and horizontal asymptotes, all holes, and all x- and y-intercepts.

OpenStudy (anonymous):

@eliassaab lol i already punched it into my phones wold ram alpha app i just don't understand the other questions what exactly is a hole?

OpenStudy (anonymous):

A hoe should be a point where the function is not continuous

OpenStudy (the_fizicx99):

0.0 that sounded wrong xD

OpenStudy (anonymous):

@tHe_FiZiCx99 lmfao ikr?

OpenStudy (anonymous):

Factor the denominator \[ x^2-2 x-3=(x-3) (x+1) \] at x=3 and x =-1 the function is not continuous and it has vertical asymptotes

OpenStudy (the_fizicx99):

She didn't mean it that way though lol xD I'm sure it was a typo.

OpenStudy (anonymous):

Also \[ \lim_{x\to \infty } \, \frac{3 x^2-12}{x^2-2 x-3}=3 \] Hence y=3 is a horizontal asymptote